OpenStudy (-welp-):

How do I find the solution set? x-2/x+4 = 2- 4/x

OpenStudy (-welp-):

\[\frac{ x-2 }{ x+4 } = 2 - \frac{ 4 }{ x }\]

OpenStudy (4815162342):

cancel x

OpenStudy (4815162342):

answer is 1.6

OpenStudy (-welp-):

Oh

OpenStudy (4815162342):

was it right?